# thanks: Corresponding author

### v2

@@ Appendix A Received intensities of Model-II and Model-III
In this appendix, we briefly present the received intensity profiles for selected polar angles $\eta$ in optically thin accretion disk Models II and III, shown in Figs. 1417, at inclination angles of $\pi/4$ and $3617\pi/$ 36.
@@
Figure 15: For Model-II at $3617\pi/$ 36 inclination, the counter-side and co-side intensities are shown for varying polar angles $\eta$. The upper and lower rows show $I_{\text{obs1}}$ and $I_{\text{obs2}}$ respectively.
@@
Figure 17: For Model-III at $3617\pi/$ 36 inclination, the counter-side and co-side intensities are shown for varying polar angles $\eta$. The upper and lower rows show $I_{\text{obs1}}$ and $I_{\text{obs2}}$ respectively.

### v3

@@ Appendix A Received intensities of Model-II and Model-III
In this appendix, we briefly present the received intensity profiles for selected polar angles $\eta$ in optically thin accretion disk Models II and III, shown in Figs. 1518, at inclination angles of $\pi/4$ and $17\pi/36$.
@@
Figure 16: For Model-II at $17\pi/36$ inclination, the counter-side and co-side intensities are shown for varying polar angles $\eta$. The upper and lower rows show $I_{\text{obs1}}$ and $I_{\text{obs2}}$ respectively.
@@
Figure 18: For Model-III at $17\pi/36$ inclination, the counter-side and co-side intensities are shown for varying polar angles $\eta$. The upper and lower rows show $I_{\text{obs1}}$ and $I_{\text{obs2}}$ respectively.